32. After aging the accounts receivable, it is estimated thaf he accou allowance account has an existing credit balance of $A net receivables would be is contra to W ng credit t us estimated that $1,200 will not be collected aU If the accounts receivable total $130,000, the O, to write off an account that has been determined to be uncollectible, the entry would include 34. After aging the accounts account has an existing debit balance of $230. The adjusting entry under the aging approach would be receivable, it is estimated that $790 will not be collected and the allowance for the amount of 35. After aging the accounts receivable, it is estimated that $420 will not be collected and the allowance account has an existing debit balance of $100. If accounts receivable is $145,000, the net receivables would be 36. Accountants argue that which of the following approaches to estimating Allowance for Bad Debt provides a realistic estimate of the net receivables? ceivable, it is estimated that $700 will not be collected and the allowance 37. After aging the accounts re account has an existing credit balance of $100. The adjusting entry for the amount of 38. The direct write-off method is method, when an account receivable is written off in one accounting he following accounting period, which of the following would be included in an acco 39 Under the direct write-off d and is collected in the period, whi erio the journal entry?

the entry would include After aging the accounts receivable, it is estimated that $790 will not be collected and the allowance account has an existing debit balance of $230. for the amount of The adjusting entry under the aging approach would be 35. After aging the accounts receivable, it is estimated that $42 0 will not be collected and the allowance account has an existing debit balance of $100. If accounts receivable is $145,000, the net receivables would be 36. Accountants argue that which of the following approaches to estimating Allowance for Bad Debts provides a realistic estimate of the net receivables? 37. After aging the accounts receivable, it is estimated that $700 will not be collected and the allowance account has an existing credit balance of $100. The adjustin nder
Answer #1

32)130,000 -1,200 = 143,800

33) Allowance for bad debts Dr.

To Accounts Receivable

34) 790 +230 =1,020

35) 145,000 -420 =143,800

36) Aging the Accounts Receivable

37)700 -100 = 600

38) Required for Income tax Purpose

39) Accounts Receivable

To Bad Debts Expense